# Always-Sparse Training by Growing Connections with Guided Stochastic Exploration

### v1

@@ 3.2 Dynamic sparse training
Our complete dynamic sparse training algorithm starts by initializing each layer as a random bipartite graph using the Erdős–Rényi random graph generation algorithm (Erdős and Rényi, 1959). The number of active connections of a layer is given by where is a parameter that controls the sparsity of the model (Mocanu et al., 2018). This initialization ensures that the number of active connections in a layer is proportional to its width, that is, . Moreover, Evci et al. (2020) showed that the Erdős–Rényi initialization also achieves better accuracy than uniform sparsity assignment. We verify their findings in Appendix F. The weights are initialized using the procedure proposed by Evci et al. (2019) which adapts standard weight initialization to take into account the actual connectivity of each unit and scales the weight distribution for each unit accordingly.
@@
Once the random sparse model is initialized, the weights of the active connections are optimized using stochastic gradient descent (SGD). Every training steps new connections are grown and the same number of connections are pruned. We provide pseudocode for our pruning and growing procedure in Algorithm 1, which operates on a single layer for simplicity. The maximum fraction of pruned active connections is cosine annealed from at the start of training to zero at . This ensures that there is more exploration in the early stages of training.
@@
To determine the set of connections to grow , we start by evaluating one of the probability distributions described in Section 3.1. Then, connections are sampled from the input and output distributions, where is a hyperparameter that adjusts the number of samples. The subset contains all randomly sampled connections that are not in the active set . Now that the subset of inactive connections is determined, their gradient magnitude is computed. The grow set contains the largest gradient magnitude connections from the subset, where is the minimum between and to ensure that the grow set is not larger than the subset. The largest elements can be found in time using the introspective selection algorithm (Musser, 1997).
@@
The pruned set , in turn, contains the active connections with the smallest weight magnitude, also determined in time. Although magnitude pruning has been observed to cause layer collapse among pruning before training methods, Tanaka et al. (2020) investigated layer collapse and found that the training process is implicitly regularizing the weight magnitudes which minimizes the risk of layer collapse for DST methods like ours. Lastly, the active set and the weights are updated to reflect the changed connections. Pseudocode for the entire training process is provided in Appendix G. It is important to stress that the dense model is not materialized at any point, only the active connections are represented, and all the operations are always sparse.
4 Experiments
At this point, in order to verify our claims, we need to determine what the minimum number of subset samples is that maintains the maximum accuracy. We answer this question in Section 4.2 and then compare the accuracy of our method against other sparsification methods in Section 4.3. We present accuracy results for ImageNet in Section 4.4. In Section 4.5, we explore the relation between model scale and accuracy for a fixed number of active connections by increasing the sparsity of larger models. Lastly, in Section 4.6, we compare the FLOPs of GSE with those of RigL.
4.1 Experiment setup
@@
We evaluate our method and compare it to baselines on the CIFAR-10 and CIFAR-100 datasets (Krizhevsky et al., 2009), in addition to the ILSVRC-2012 ImageNet dataset (Deng et al., 2009). To establish how each method compares across model architectures we experiment with two convolutional neural networks (CNNs), ResNet-56 (He et al., 2016) and a 4 times downscaled version of VGG-16 (Simonyan and Zisserman, 2014), in addition to ViT (Dosovitskiy et al., 2020) modified for small datasets (Beyer et al., 2022) such that all models have roughly the same number of parameters. The bias and normalization parameters are kept dense since they only contribute marginally to the size and cost of a model. Similar to Wang et al. (2019), we evaluate all methods at 90%, 95%, and 98% sparsity. The high levels of sparsity emphasize the differences between the methods. In addition, in Appendix L, we show that sparse matrix formats for unstructured sparsity become more efficient for matrix multiplication from 90% sparsity compared to dense matrix formats, further motivating the need for such high levels of sparsity. We repeat each experiment three times with random initializations and report the mean and plot the 95th percentile. We adopt the implementation of the baselines based on their available code.
@@
All experiments use the same optimization settings in order to isolate the differences in sparsification. Similar to Evci et al. (2020) and Lubana and Dick (2020), we use SGD with a momentum coefficient of 0.9, an regularization coefficient of 0.0001, and an initial learning rate of 0.1 which is dropped by a factor of 10 at epochs 80 and 120. The ViT experiments have one exception, they use an initial learning rate of 0.01. We use a batch size of 128 and train for 200 epochs. We also apply standard data augmentation to the training data, described in Appendix D. Running all the reported experiments took on the order of 2000 GPU-hours with a single NVIDIA Tesla V100 16 GB GPU per training run on an internal cluster.
@@
The LTH procedure proposed by Frankle and Carbin (2018), denoted as Lottery in the results, uses iterative pruning with iterations at 0%, 50%, 75%, 90%, 95% and 98% sparsity. The gradual pruning method by Zhu and Gupta (2017), denoted as Gradual in the results, reaches the target sparsity at the second learning rate drop and prunes connections every 1000 steps. Following Tanaka et al. (2020), the pruning before training methods use a number of training samples equal to ten times the number of classes to prune the dense model to the target sparsity. Except for SynFlow, which uses 100 iterations with an exponential decaying sparsity schedule to prune the dense model. All the DST methods use the same update schedule: the connections are updated every 1000 steps, and similar to Evci et al. (2020), the fraction of pruned active connections is cosine annealed from 0.2 to 0.0 at the second learning rate drop. This is because Liu et al. (2021a) showed that DST methods struggle to recover from pruning when the learning rate is low. The pruning and growing procedures are applied globally for all methods, maintaining the overall sparsity. This approach generally outperforms local, layer-wise sparsity (Mostafa and Wang, 2019).
4.2 Effect of subset sample size
@@
First, we want to determine how the size of the subset affects the accuracy. In addition, we are interested in comparing the accuracy obtained by the three distributions: uniform, gradient upper bound (GraBo), and gradient estimate (GraEst). The number of subset samples (bounding the subset size) is set proportional to the number of active connections , with and . The results are shown in the first three columns of Figure 2 and include the baseline accuracy of RigL for comparison.
@@
Interestingly, uniform sampling demonstrates strong performance while we hypothesized that the biased distributions would be beneficial for training. This shows that exploring connections which, at first, do not seem beneficial for training, can in fact be quite helpful. It would be interesting for future research to investigate exactly which training dynamics are enabled by uniform sampling. Since uniform sampling is also the most efficient distribution, we use it as our method in the following experiments.
4.3 Comparison with related work
We compare our method to a broad range of sparsification methods in Table 1, the best accuracy among the sparse training methods is bolded. The Static method denotes training a static random sparse model. The other baselines are Lottery (Frankle and Carbin, 2018), Gradual (Zhu and Gupta, 2017), SNIP (Lee et al., 2018), GraSP (Wang et al., 2019), SynFlow (Tanaka et al., 2020), SET (Mocanu et al., 2018), RigL (Evci et al., 2020), and Top-KAST (Jayakumar et al., 2020). Our results use GSE with uniform sampling and .
@@ 4.3 Comparison with related work
Among the sparse training methods (from Static downwards in Table 1) our method outperforms the other methods consistently over datasets, sparsities, and model architectures. While at 90% sparsity all the methods achieve comparable accuracy, at the extreme sparsity rate of 98%, differences between methods become more evident. At 98% sparsity our method outperforms all the before training sparsification methods by a significant margin, 4.7% on average. GSE also consistently matches or surpasses both RigL and Top-KAST, while being more efficient. For example, with VGG-16 on CIFAR-10, GSE achieves 1.4% and 4.6% higher accuracy than RigL and Top-KAST, respectively.
Gradually pruning connections during training (Gradual) can still improve the accuracy, especially in the extreme sparsity regimes. However, this comes at the cost of training the dense model, thus requiring significantly more FLOPs than our method. This observation serves as motivation for further research into dynamic sparse training methods that improve on our accuracy while preserving our efficiency advantage. Notably, our method outperforms Lottery on the CNN models, while Lottery requires multiple rounds of training.
In Appendix H, we report the test accuracy of each method at every epoch throughout training, showing that our method trains as fast as the other sparse training methods. Additionally, in Appendix M, we provide plots of the layer-wise sparsity at the end of training for each combination of method, model, and sparsity. This information provides insight into how each method distributes the available connections over the layers of a model.
4.4 ImageNet
@@
To determine the efficacy and robustness of GSE on large models and datasets, in this third experiment, we compare the accuracy of ResNet-50 trained on ImageNet at 90% sparsity. In Table 2, the accuracy obtained with GSE is compared against the baselines from the preceding section, with the additions of DSR (Mostafa and Wang, 2019) and SNFS (Dettmers and Zettlemoyer, 2019). Since this is a common benchmark among sparse training methods, the accuracy results of the baselines were obtained from prior publications.
@@
Our experiments use the aforementioned hyperparameters with the following exceptions: we train for 100 epochs and use learning rate warm-up over the initial 5 epochs, gradually reaching a value of 0.1; subsequently, the learning rate is reduced by a factor of 10 every 30 epochs; and we use label smoothing with a coefficient of 0.1. Our results use GSE with uniform sampling and . The results in Table 2 align with those for CIFAR, and further underscore GSE’s superior performance over other sparse training methods. Note also that GSE again outperforms less efficient methods such as SNFS, RigL, and Top-KAST. These results collectively reinforce that GSE is robust and able to scale to large models and datasets.
4.5 Model scaling
@@
In this experiment, we investigate how the size of a model affects the accuracy when the number of active connections is kept constant. We scale the number of filters in convolution layers and units in fully-connected layers to obtain a wider model. The results, presented in Figure 3, show that training wider CNN models translates to significantly increased accuracy, 6.5% on average at 4 times wider. These findings support those from Zhu and Gupta (2017) and Kalchbrenner et al. (2018), however, our method enables a sparse model to be trained whose dense version would be too computationally intensive.
@@
The same trend does not hold for ViT, as its accuracy decreases with increasing scale. Appendix J reveals that the accuracy increases slower for larger ViT models, while for CNNs the accuracy increases faster as they get larger. Moreover, the accuracy of larger dense ViT models plateaus at 62.2%, as shown in Appendix K, indicating that this phenomenon is independent of our method. ViT models are known to require large amounts of training data to achieve optimal accuracy (Dosovitskiy et al., 2020), we conjecture that this contributes to the trend observed with ViT. In Appendix I, we observe that both ViT and CNN models do benefit considerably from extended training, with a 5.3% average increase in accuracy for training twice as long.
4.6 Floating-point operations
Lastly, to determine the efficiency of GSE, we compare its FLOPs to those of RigL in Figure 4. The FLOPs are obtained for ResNet-50 on ImageNet using the method described by Evci et al. (2020), with . The results show that at 99% sparsity GSE uses 11.8% fewer FLOPs than RigL, at higher sparsities GSE uses considerably fewer FLOPs. For illustration, the right plot extends the sparsity to that of various animal brains which underscores the significance of our always sparse training method, as RigL uses times more FLOPs to train a model with the same sparsity as the cat brain.
4.7 Limitations
While we experiment on a range of sparsities, datasets, models, and baseline methods, we acknowledge that further exploration would enhance the robustness of our findings. Following previous work, our experiments focus on the domain of computer vision, and although we demonstrate the effectiveness of our approach within this domain, it remains to be seen how our results generalize to other domains such as natural language processing and reinforcement learning, among others. Despite this, we believe that our proposed always-sparse training algorithm, supported by its linear time complexity with respect to the model width during training and inference, represents a significant advancement. Furthermore, our method outperforms existing sparse training algorithms in terms of accuracy, as demonstrated on the CIFAR and ImageNet datasets using ResNet, VGG, and ViT models, alongside a comprehensive comparison with a range of sparsification techniques.
Secondly, our accuracy results were obtained by simulating sparse networks using binary masks. This is standard practice among sparse training methods because support for sparsity in machine learning frameworks is still in the early days, making it challenging to train models using sparse matrix formats for the weights. We note that there is nothing about our method that prevents it from being accelerated on GPUs, yet the significant engineering required to develop an optimized GPU kernel falls outside the scope of this work. To ensure that our claims are still validated, we compare the FLOPs of our method with those of RigL in Section 4.6. Moreover, we present an unstructured sparse matrix multiplication benchmark in Appendix L, which shows that a 10 times reduction in execution time is achieved at 99% sparsity, compared to dense matrix multiplication. Thus, evidently showing that real speedups can be achieved, even with unstructured sparsity.
5 Conclusion
We presented an always-sparse training algorithm using guided stochastic exploration of the connections which has a linear time complexity with respect to the model width and improves upon the accuracy of previous sparse training methods. This is achieved by only evaluating gradients for a subset of randomly sampled connections when changing the connections. The three distributions for sampling the subset—uniform, an upper bound of the gradient magnitude, and an estimate of the gradient magnitude—all showed similar trends. However, uniform consistently achieves among the highest accuracy and is also the most efficient. We therefore conclude that the uniform distribution is the most appropriate for sparse training. Interestingly, we showed that training larger and sparser CNN models results in higher accuracy for the same number of active connections. Lastly, the number of floating-point operations of our method decrease considerably faster than those for RigL as the sparsity increases. This observation underscores the efficiency and scalability of our method, which is promising for training increasingly larger and sparser ANNs.
Appendix A Connection sampling distributions
In this section we provide a more detailed discussion on the distributions which were considered for sampling the subset of inactive connections.
A.1 Gradient magnitude upper bound (GraBo)
The gradient magnitude for a single connection can be expressed as follows:
where and are the activations and gradients of the loss at the output units of the -th layer, respectively. When the batch size is one, the sum disappears, simplifying the gradient magnitude to a rank one matrix which can be used to sample connections efficiently, inducing a joint probability distribution that is proportional to the gradient magnitude only when the batch size is one. In practice, however, training samples come in mini batches to reduce the variance of the gradient. Therefore, we experiment with sampling the connections proportionally to the following upper bound of the gradient magnitude instead, which enables efficient sampling, even with mini batches:
The proof for Equation 2 uses the triangle inequality and is provided in Appendix B. This upper bound has the property that it becomes an equality when . Connections are then sampled from the probability distributions in Equation 3. The implementation of this distribution does not require any modifications to the model, and only needs a single forward and backward pass to evaluate.
A.2 Gradient magnitude estimate (GraEst)
The reason we use a bound on the absolute gradient is that in general the sum of products cannot be written as the product of sums, that is, . However, it is possible to satisfy this equality in expectation. This was shown by Alon et al. (1996) who used it to estimate the second frequency moment, or the norm of a vector. Their work is called the AMS sketch, AMS is an initialism derived from the last names of the authors. They then showed that this also works in the more general case to estimate inner products between vectors (Alon et al., 1999). This is achieved by assigning each index of the vector a random sign, i.e. , resulting in the following equality:
We can use their finding to construct probability distributions and in Equation 5 such that their induced joint probability distribution is proportional to the gradient magnitude in expectation. This distribution can also be computed in a single forward and backward pass and requires no modifications to the model, apart from introducing a mechanism to sample the random signs .
Appendix B Gradient magnitude upper bound proof
Lemma B.1.
The gradient magnitude of the loss with respect to the parameters of a fully-connected layer has the following upper bound:
where , , and are the weight matrix, activations and gradients of the loss at the output units of the -th layer, respectively. The number of units in the -th layer is denoted by , and is the batch size.
Proof.
The proof starts with the definition of the gradient of the fully-connected layer, which is calculated during back propagation as the activations of the previous layer multiplied with the back propagated gradient of the output units of the layer. We then write this definition for a single connection in the weight matrix and use the triangle inequality to specify the first upper bound. In Equation 8 we rewrite the upper bound of Equation 7 as all the cross products between the batch dimensions minus all the non-identical cross terms. Note that all the non-identical cross terms are positive because of the absolute, therefore we get another upper bound by removing the non-identical cross terms which we then write in matrix notation in Equation 9.
Appendix C Data normalization
We normalize all the training and test data to have zero mean and a standard deviation of one. The dataset statistics that we used are specified below, where is the mean and the standard deviation. The values correspond to the red, green, and blue color channels of the image, respectively.
CIFAR-10
CIFAR-100
ImageNet
Appendix D Data augmentation
We applied standard data augmentation as part of our training data pipeline. The specific augmentation techniques are specified below. We apply the data normalization described in Appendix C after the augmentation transformations.
CIFAR-10
We pad the image with 4 black pixels on all sides and randomly crop a square of 32 by 32 pixels. We then perform a random horizontal flip of the crop with a probability of 0.5.
CIFAR-100
We pad the image with 4 black pixels on all sides and randomly crop a square of 32 by 32 pixels. We then perform a random horizontal flip of the crop with a probability of 0.5.
ImageNet
We randomly crop a square of 224 by 224 pixels. We then perform a random horizontal flip of the crop with a probability of 0.5. For the test set, we resize the images to be 256 pixels followed by a square center crop of 224 pixels.
Appendix E Convolutional layer subset sampling
In this section we describe how our growing and pruning algorithm can be applied to convolutional layers. While a weight in a fully-connected layer is specified by the input and output units that it connects, a weight in a 2D convolutional layer is specified by the input channel, its and coordinate on the filter, and the output channel. It thus seems that it would require four discrete probability distributions to sample a weight of a 2D convolutional layer, instead of two for fully-connected layers. However, one can alternatively view a convolutional layer as applying a fully-connected layer on multiple subsections (patches) of the input, flattening the input channels and filters makes the weights of a convolutional layer identical to a fully-connected layer. The multiple subsections of the input can then simply be treated as additional batches. With this realization all the methods that we explain throughout the paper for fully-connected layers directly apply to convolutional layers.
Appendix F Sparse graph initialization comparison
In Figure 5, we plot the accuracy obtained by assigning the sparsity uniformly over all the layers compared to using the Erdős–Rényi initialization for SET, RigL, and our method. Note that Ours Uniform (Erdős–Rényi) denotes uniform sampling of the subset connections and the Erdős–Rényi initialization. The Erdős–Rényi initialization obtains better accuracy while ensuring that the number of active connections per layer is proportional to the layer width.
Appendix G Pseudocode
We provide pseudocode for our complete dynamic sparse training process in the case of the GraBo or GraEst distributions in Algorithm 2. The uniform distribution is simpler as it does not require any aggregation of activity. For simplification, the pseudocode assumes that the prune-grow step only uses a single batch of training samples, allowing the aggregation for the probability distributions f and g to be computed in the same pass as the generation of the various connection sets. This is also the scenario we tested with in all our experiments.
Appendix H Training progress
In Figure 6, we provide plots for the test accuracy during training for each baseline method and our method. It can be seen that our method consistently achieves among the highest test accuracy throughout training compared to the sparse training methods. The training progress of Gradual stands out because its accuracy increases the most at first, this is because it starts out as a dense model while it gradually increases the model sparsity, as a result, the accuracy goes down towards the second learning rate drop when the target sparsity is reached.
Appendix I Extended training

### v2

@@ III-B Dynamic sparse training
Our complete dynamic sparse training algorithm starts by initializing each layer as a random bipartite graph using the Erdős–Rényi random graph generation algorithm [64]. The number of active connections of the -th layer is given by , where is a parameter that controls the sparsity of the model [33]. This initialization ensures that the number of active connections in a layer is proportional to its width, that is, . Moreover, the Erdős–Rényi initialization also achieves better accuracy than uniform sparsity assignment [35], as verified in Appendix -F. Once the initial connections are sampled, the weights are initialized so that the activations are normally distributed [65].
@@
During training, the weights of the active connections are optimized using stochastic gradient descent (SGD). Every training steps new connections are grown, which are added with a weight of zero, and the same number of connections are pruned. We provide pseudocode for our pruning and growing procedure in Algorithm 1, which operates on a single layer for simplicity. The maximum fraction of pruned active connections is cosine annealed from at the start of training to zero at . This ensures that there is more exploration in the early stages of training.
@@
To determine the set of connections to grow , we start by evaluating one of the probability distributions described in Section III-A. Then, connections are sampled from the input and output distributions, where is a hyperparameter that adjusts the number of samples. The subset contains all randomly sampled connections that are not in the active set . Now that the subset of inactive connections is determined, their gradient magnitude is computed. The grow set contains the largest gradient magnitude connections from the subset, where is the minimum between and to ensure that the grow set is not larger than the subset. The largest elements can be found in time using the introspective selection algorithm [66].
@@
The pruned set , in turn, contains the active connections with the smallest weight magnitude, also determined in time. Although magnitude pruning has been observed to cause layer collapse among pruning before training methods, the training process implicitly regularizes the weight magnitudes which minimizes the risk of layer collapse for DST methods like ours [32]. Lastly, the active set and the weights are updated to reflect the changed connections. Pseudocode for the entire training process is provided in Appendix -G. It is important to stress that the dense model is not materialized at any point, only the active connections are represented, and all the operations are always sparse.
IV Experiments
@@
We evaluated our method and compared it to baselines on the CIFAR-10 and CIFAR-100 datasets [67], in addition to the ILSVRC-2012 ImageNet dataset [68]. To establish how each method compares with different model architectures, we experimented with two convolutional neural networks (CNNs), ResNet-56 [69] and a 4 times downscaled version of VGG-16 [70], in addition to ViT [71] modified for small datasets [72] so that all models have roughly the same number of parameters. The bias and normalization parameters are kept dense since they only contribute marginally to the size and cost of a model. Like Wang et al. [41], we evaluated all methods at 90%, 95%, and 98% sparsity. The high levels of sparsity emphasize the differences between the methods. In addition, in Appendix -K, we show that sparse matrix formats for unstructured sparsity become more efficient for matrix multiplication from 90% sparsity compared to dense matrix formats, further motivating the need for such high levels of sparsity. We repeat each experiment three times with random initializations. The plots report the mean and the 95th percentile, while the tables report the mean and standard deviation. We adopt the implementation of the baselines based on their available code.
@@
All experiments use the same optimization settings to isolate the differences in sparsification. Following Evci et al. [35] and Lubana et al. [57], we use SGD with a momentum coefficient of 0.9, an regularization coefficient of 0.0001, and an initial learning rate of 0.1 which is dropped by a factor of 10 at epochs 80 and 120. The ViT experiments have one exception: they use an initial learning rate of 0.01. We use a batch size of 128 and train for 200 epochs. We also apply standard data augmentation to the training data, described in Appendix -D.
@@
The LTH procedure proposed by Frankle et al. [29], denoted as Lottery in the results, uses iterative pruning with iterations at 0%, 50%, 75%, 90%, 95% and 98% sparsity. The gradual pruning method by Zhu et al. [22], denoted as Gradual in the results, reaches the target sparsity at the second learning rate drop and prunes connections every 1000 steps. Following Tanaka et al. [32], the pruning before training methods use a number of training samples equal to ten times the number of classes to prune the dense model to the target sparsity. Except for SynFlow, which uses 100 iterations with an exponential decaying sparsity schedule to prune the dense model. All DST methods use the same update schedule: the connections are updated every 1000 steps, and similar to Evci et al. [35], the fraction of pruned active connections is cosine annealed from 0.2 to 0.0 at the second learning rate drop. This is because Liu et al. [50] showed that DST methods struggle to recover from pruning when the learning rate is low. The pruning and growing procedures are applied globally for all methods, maintaining the overall sparsity. This approach generally outperforms local, layer-wise sparsity [34].
IV-A Effect of subset sample size
@@
First, we want to determine how the size of the subset affects the accuracy. In addition, we are interested in comparing the accuracy obtained by the three distributions: uniform, gradient upper bound (GraBo), and gradient estimate (GraEst). The number of subset samples (which is a bound on the subset size) is set proportional to the number of active connections , with and . The results are shown in the first three columns of Figure 3 and include the baseline accuracy of RigL for comparison.
@@
Surprisingly, uniform sampling demonstrates strong performance while we hypothesized that biased distributions would be beneficial for training. This shows that exploring connections which, at first, do not seem beneficial for training, can in fact be quite helpful. It would be interesting for future research to investigate exactly which training dynamics are enabled by uniform sampling. Since uniform sampling is also the most efficient distribution, we use it as our method in the following experiments.
IV-B Comparison with related work
We compare our method with a wide range of sparsification methods in Table I on CIFAR-10 and CIFAR-100. The best accuracy among the sparse training methods is in bold. The Static method denotes training a static random sparse model. The other baselines are Lottery [29], Gradual [22], SNIP [31], GraSP [41], SynFlow [32], SET [33], RigL [35], and Top-KAST [36]. Our results use GSE with uniform sampling and .
Among the sparse training methods (from Static downwards in Table I) our method outperforms the other methods consistently over datasets, sparsities, and model architectures. While at 90% sparsity all methods achieve comparable accuracy, at the extreme sparsity rate of 98%, differences between methods become more evident. At 98% sparsity our method outperforms all the before training sparsification methods by a significant margin, 4.7% on average. GSE also consistently matches or exceeds both RigL and Top-KAST, while being strictly more efficient. For example, with VGG-16 on CIFAR-10, GSE achieves 1.4% and 4.6% higher accuracy than RigL and Top-KAST, respectively.
Gradually pruning connections during training (Gradual) can still improve the accuracy, especially in the extreme sparsity regimes. However, this comes at the cost of training the dense model, thus requiring significantly more FLOPs than our method. This observation serves as motivation for further research into dynamic sparse training methods that improve our accuracy while preserving our efficiency advantage. Notably, our method outperforms Lottery on the CNN models, while Lottery requires multiple rounds of training. In Appendix -H, we report the test accuracy of each method at every epoch throughout training, showing that our method trains as fast as the other sparse training methods.
IV-C ImageNet
To determine the efficacy and robustness of GSE on large models and datasets, in this third experiment, we compare the accuracy of ResNet-50 trained on ImageNet at 90% sparsity. In Table II, the accuracy obtained with GSE is compared against the baselines from the preceding section, with the additions of DSR [34] and SNFS [61]. Since this is a common benchmark among sparse training methods, the accuracy results of the baselines were obtained from prior publications.
@@ IV-C ImageNet
Our experiments use the aforementioned hyperparameters with the following exceptions: we train for 100 epochs and use learning rate warm-up over the initial 5 epochs, gradually reaching a value of 0.1; subsequently, the learning rate is reduced by a factor of 10 every 30 epochs; and we use label smoothing with a coefficient of 0.1. Our results use GSE with uniform sampling and . The results in Table II show a 0.2% improvement in model accuracy, which is the largest increase since the introduction of SNFS [61] in 2019 and sets a new state-of-the-art in dynamic sparse training. These results align with those for CIFAR, and further underscore GSE’s performance over other sparse training methods. Note that GSE not only obtains a higher accuracy, it is also more efficient than methods such as SNFS, RigL, and Top-KAST. These results collectively reinforce GSE’s robustness and its ability to scale to large models and datasets.
IV-D Model scaling
@@
In this experiment, we investigate how the size of a model affects the accuracy when the number of active connections is kept constant. We scaled the number of filters in convolution layers and units in feedforward layers to obtain a wider model. The results, presented in Table III, show that training wider CNN models translates to significantly higher accuracy, 6.5% on average at 4 times wider. These findings support those of Zhu et al. [22] and Kalchbrenner et al. [23], however, our method enables a sparse model to be trained whose dense version would be too computationally intensive.
| Width / Sparsity | 1x / 0% | 2x / 75% | 3x / 89% | 4x / 94% |
| ResNet-56 | 70.1±0.4 | 72.3±0.4 | 73.8±0.6 | 74.2±0.3 |
| VGG-16 | 62.3±0.8 | 68.4±0.7 | 70.4±0.8 | 71.2±0.3 |
| ViT | 58.8±0.5 | 58.3±0.2 | 57.3±0.7 | 56.6±0.1 |
@@
The same trend does not hold for ViT, as its accuracy decreases with increasing scale. Appendix -I reveals that the accuracy increases slower for larger ViT models, while for CNNs the accuracy increases faster as they get larger. Moreover, the accuracy of larger dense ViT models plateaus at 62.2%, as shown in Appendix -J, indicating that this phenomenon is independent of our method. ViT models are known to require large amounts of training data to achieve optimal accuracy [71], we conjecture that this contributes to the trend observed with ViT.
IV-E Extended training